OpenStudy (anonymous):

I do not understand how to solve this problem. Can someone show me how to do it step by step? Q= An investor invested a total of $3,100 in two mutual funds. One earned 4% profit while the other earned a 2% profit. If the investors total profit was $100, how much was invested in each fund? 4% was= 2% was=

OpenStudy (anonymous):

call the amount the investor put into the first mutual fund 'x' and the amount invested in the second 'y'. We should have the following relations: x+y=3100 1.02x+1.04y=100

OpenStudy (anonymous):

You should be able to solve it from here since we have two equations and two unknowns...let me know if you get stuck.

OpenStudy (anonymous):

I will attempt to solve it now.

OpenStudy (anonymous):

actually I'm thinking I made a mistake on the second equation. I think it should be: 0.02x+0.04y=100 since we only want the profit

OpenStudy (anonymous):

0.02x+0.04(3100-x)=100 -0.02x=-24 x=$1200 so, y=3100-1200=$1900

OpenStudy (anonymous):

4%=1900 2%==1200

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

Thank you, I have been trying to figure that out for the past 30 mins.

OpenStudy (anonymous):

no prob :)